# The role of gaps in digitized counterdiabatic QAOA for fully-connected spin models

### v1

@@ I Introduction
Many QAOA variants have proliferated over the years to further improve the performance of the original algorithm [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 27, 16, 33, 34, 35]. A non-exhaustive list includes: QAOA$+$[36], which improves the conventional QAOA adding an additional problem-independent layer with multiple parameters; adaptive-bias QAOA [26], introducing local fields to QAOA operators; adaptive QAOA [27], a QAOA variant iteratively selecting mixers based on a systematic gradient criterion; and recursive-QAOA [28], aiming to decrease problem size by eliminating unnecessary qubits through a non-local scheme. A promising variant is the digitized counterdiabatic QAOA (QAOA-CD) [33, 34, 35], which incorporates an additional driving Hamiltonian inspired by quantum shortcuts to adiabaticity [37, 38, 39]. This additional Hamiltonian leads to faster convergence to the ground state energy, allowing for a reduction in the circuit depth.
In this paper, we apply the variant of QAOA originally developed in Ref. [39] to study the randomly weighted MaxCut problem on 2-regular graphs, to explore its suitability to to fully-connected graphs in which long-range all to all interactions significantly amplify the computational complexity with respect to local interactions. In this setting, our goal is twofold. On the one hand, we will show that the counterdiabatic variants of QAOA are also feasible for fully-connected models. On the other hand, we will show through an in-depth statistical study that the performance of QAOA and its variants depend on the spectral properties of the target Hamiltonian. We can in fact distinguish “easy” and “hard” Hamiltonians depending on the value of the gap between the target ground state and the first excited states, in a similar fashion to adiabatic quantum computing (AQC). In AQC, the minimum instantaneous gap between the time-dependent ground state and first excited state classifies the hardness of the optimization problem as it determines the minimum time scale necessary for successful adiabatic state preparation [40]. In QAOA, instead, where continuous evolutions are replaced by digital steps and it is not possible to identify a continuous time-dependent generator of the quantum circuit, this role is played by the gap of the target Hamiltonian.
@@ II.1 Fully-connected spin model
where $\sigma_{i}^{Z},\sigma_{j}^{Z}$ are the Pauli matrices acting on respectively the $i^{\text{th}}$ and $j^{\text{th}}$ spins and $J_{ij}=\mathcal{U}([-1,1])$ are random couplings distributed according to the uniform distribution $\mathcal{U}$ in the given range. The cost function of QAOA, as discussed in Sec. III, is given by the expectation value of $H_{T}$ on a parametric quantum circuit. As a proof of principle we focus on random instances with a small number of qubits, $N=5$, and analyze $n=600$ samples (which gives an accurate sampling of the coupling distribution without requiring unfeasible computational resources) of the disordered interactions, each of them with different couplings $J_{ij}$. In the case $J_{ij}=1$ finding the ground state of the Hamiltonian of Eq. (2) is the same as finding the maximum of the cost function $f$ defined in Eq. (1). Due to the $Z_{2}$ symmetry of the Hamiltonian of Eq. (2) all the eigenvalues are doubly degenerate.
@@ III QAOA and variants
QAOA is a hybrid quantum-classical algorithm proposed by Farhi et al [9]. It consists of a quantum part and a classical part. The objective of the algorithm is to find the ground state of a target Hamiltonian $H_{T}$. The algorithm works as follows. We consider two Hamiltonians: $H_{T}$, representing our target Hamiltonian whose ground state solves our optimization problem, and $H_{X}$, a simple Hamiltonian easy to handle, encoding quantum fluctuations:
@@
|$\ket{\psi_{\text{2CD}}^{(p)}(\vec{\gamma},\vec{\beta},\vec{\alpha},\vec{\delta},\vec{\zeta})}=$
@@ IV Results
In this section, we discuss results achieved by studying the disordered fully-connected spin model of Eq. (2) within the QAOA, QAOA-CD, and QAOA-2CD frameworks. Our results are averaged over a set of $n=600$ random instances. Disorder-averaged quantities are calculated as
@@
The fidelity of Eq. (16) as well as the residual energy of Eq. (12) are used in the following to assess the performances of the algorithm and eventually connect them with the spectral properties of $H_{T}$. In quantum annealing [40, 53, 64], the minimal gap during the annealing dynamics determines the fidelity of the algorithm due to the occurrence of Landau-Zener transitions. In QAOA such strict relation between the gaps and the fidelity is not known nor well established, however a dense spectrum in the vicinity of the ground state, makes the system more prone to be “captured” by local minima during the hybrid optimization procedure [65]. To highlight the role of the gaps in the QAOA, in the following we will relate the QAOA performances for various instances keeping track of the gap $\Delta_{eg}$ between the ground and the first excited state.
@@
Figure 2: (a) Residual energy versus QAOA step. Grey lines represent $\varepsilon_{\mathrm{res,}i}$ of all the $n=600$ random instances and the red line their average value $\braket{\varepsilon_{\mathrm{res}}}$. (b-d) Residual energy $\varepsilon_{\mathrm{res,}i}$ as a function of the minimal gap $\Delta_{eg}$ for QAOA at steps (b) $p=1$, (c) $p=2$ and (d) $p=7$. Three distinct regions based on the energy gap separating the ground state from the first excited state can be singled out for each instance: I-zone (green circles), corresponding to instances with $\Delta_{eg}\in\left[0,0.44\right]$, II-zone (blue triangles) for instances with $\Delta_{eg}\in\left[0.44,1.02\right]$, III-zone (red squares) for instances with $\Delta_{eg}\in\left[1.02,4.74\right]$. The horizontal line within each sector depicts the corresponding average value of that sector. We see that for $p=1$ these average values grow with increasing of $\Delta_{eg}$. This trend starts to change from step $p=2$ (panel c), whereas III-zone (panel d) shows a trend inversion (average value decreasing with the minimal gap).
@@
Figure 3: (a) Fidelity versus QAOA step. Grey lines represent values $F_{i}$ of all $n=600$ random instances and the red line their average value $\braket{F}$. (b-d) Fidelity $F_{i}$ as a function of the minimal gap $\Delta_{eg}$ for QAOA at step (b) $p=1$, (c) $p=2$ and (d) $p=7$. In panels (b-d) the three distinct regions described in caption of Fig. 2 are evidenced. The horizontal line within each sector depicts the corresponding average value of that sector. For each step $p$ the fidelity closest to one is obtained for instances with $\Delta_{eg}$ belonging to III-zone.
@@ IV.1 QAOA
In Fig. 2(a) we show the residual energy as a function of the number $p$ of iterations of QAOA. Each grey line corresponds to the residual energy $\varepsilon_{\mathrm{res,}i}$ for a random instance, while the red bold curve indicates the average residual energy $\braket{\varepsilon_{\mathrm{res}}}$ over all the $n$ instances. The average residual energy at the step $p=10$ is of the order of $\braket{\varepsilon_{\mathrm{res}}^{(10)}}\simeq(2\pm 3)\cdot 10^{-3}$, consistent with a zero residual energy at large $p$ which means that, on average, QAOA gives a good approximation of the exact solution at $p=10$. However the finite standard deviation indicates that some of the instances have yet to converge to their respective ground states, justifying the need for counterdiabatic corrections at finite $p$.
Is it possible to identify a common pattern among the instances that converge more slowly? In order to answer this question, we investigate the connection between the gap $\Delta_{eg}$ and complexity of the algorithm, following what is done in Ref. [66] for quantum annealing. In Figs. 2(b–d), we show scatter plots of the residual energy of each instance as a function of its gap $\Delta_{eg}$, at some fixed steps $p$. In particular, we divide the scatter plots in three regions containing the same number (200) of instances: in I there are problems with $\Delta_{eg}\in\left[0,0.44\right]$, in II $\Delta_{eg}\in\left[0.44,1.02\right]$ and in III $\Delta_{eg}\in\left[1.02,4.74\right]$. In panel (b) we show the scatter plots for $p=1$, in (c) $p=2$ and in (d) $p=7$. In panel (b), with $p=1$, we see that the larger is the gap the bigger is the residual energy; this behavior is confirmed by the average value of residual energy in each region. Starting from step $p=2$ we note that this trend reverses and becomes more and more marked as the number of steps $p$ increases, as we see comparing panels (c) and (d). In panel (c) the average residual energies of the regions II and III have similar values, then in the panel (d), for $p=7$, the average in the region III is lower than that in region I. The numerical values of the average residual energies in each region can be found in Appendix A.
@@
However, the correlation between $\Delta_{eg}$ and the values of residual energy and fidelity in QAOA could be related to the fact that in the presence of a dense spectrum of the Hamiltonian $H_{T}$ there is a large likelihood of identifying local minima in close proximity to the ground state energy.
@@ IV.2 QAOA-CD and QAOA-2CD
Figure 6: Fidelity $F_{i}$ of QAOA-2CD for all $n=$ 600 random instances versus $\Delta_{eg}$ for $p=1$ (a) and $p=2$ (b). Each point in the scatter plot represents the fidelity $F_{\mathrm{2CD,i}}$ for a single instance. Residual energy $\varepsilon_{\text{2CD,i}}$ for all $n=$ 600 different instances vs $\Delta_{eg}$ of QAOA-2CD for $p=1$ (c) and $p=2$ (d). Graphs delineate three distinct regions based on the energy gap $\Delta_{eg}$, as illustrated in caption of Fig. 2. The horizontal line within each sector depicts the corresponding average value of that sector. Fidelity in QAOA-2CD increases with $\Delta_{eg}$ as for QAOA. It is not possible to appreciate a specific behavior for residual energy.
@@
Table A2: Average residual energies $\braket{\varepsilon_{\mathrm{res}}}$, $\braket{\varepsilon_{\mathrm{res,CD}}}$ and $\braket{\varepsilon_{\mathrm{res,2CD}}}$ for QAOA, QAOA-CD and QAOA-2CD, respectively, together with respective standard deviations for all the analyzed $p$ steps. For each fixed $p$, the QAOA-2CD algorithm outperforms both the QAOA-CD and QAOA algorithms. Value denoted as $\sim 0$ have standard deviations $\leq 10^{-4}$.
@@ V Conclusions
To summarize, in this manuscript we have applied digitized counterdiabatic QAOA algorithms to study fully-connected spin models. In particolar we have compared the performances of QAOA, QAOA-CD, QAOA-2CD on a fully-connected disordered MaxCut problem with $N=5$ vertices. We analyzed $n=600$ random instances quantifying the quality of the algorithm with the residual energy (i. e., distance between the energy of the real solution and the calculated solution) and the fidelity (i. e., the overlap between the exact groundstate and the calculated state). Consistently with our prior findings [39], we observe that the QAOA-2CD overperforms the other variants, at the price of an increased number of classical parameters to optimize at fixed p.
We also analyzed the performances of the algorithms in connection with the gap between the ground state manifold and the first excited state, showing that in general, large gap instances are easier to be solved with this approach. While the role of spectral gaps are well known in QA, in QAOA their role is more subtle and this result is somehow nontrivial. The underlying, however, is that a dense spectrum in the vicinity of the ground state, makes the system more prone to be “captured” by local minima during the hybrid optimization procedure. This approach could be used as a starting point to train a neural network with datasets sorted by spectral gaps to approximate efficiently the QAOA angles as done in Ref. [66] for QA.

### v2

@@ I Introduction
Many QAOA variants have proliferated over the years to further improve the performance of the original algorithm [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 27, 16, 33, 34, 35, 36]. A non-exhaustive list includes: QAOA$+$[37], which improves the conventional QAOA adding an additional problem-independent layer with multiple parameters; adaptive-bias QAOA [26], introducing local fields to QAOA operators; adaptive QAOA [27], a QAOA variant iteratively selecting mixers based on a systematic gradient criterion; and recursive-QAOA [28], aiming to decrease problem size by eliminating unnecessary qubits through a non-local scheme. A promising variant is the digitized counterdiabatic QAOA (QAOA-CD) [33, 34, 35], which incorporates an additional driving Hamiltonian inspired by quantum shortcuts to adiabaticity [38, 39, 40]. This additional Hamiltonian leads to faster convergence to the ground state energy, allowing for a reduction in the circuit depth.
In this paper, we apply the variant of QAOA originally developed in Ref. [40] to study the randomly weighted MaxCut problem on 2-regular graphs, to explore its suitability to fully-connected graphs in which long-range all to all interactions significantly amplify the computational complexity with respect to local interactions. In this setting, our goal is twofold. On the one hand, we will show that the counterdiabatic variants of QAOA are also feasible for fully-connected models. On the other hand, we will show through an in-depth statistical study that the performance of QAOA and its variants depend on the spectral properties of the target Hamiltonian. We can in fact distinguish “easy” and “hard” Hamiltonians depending on the value of the gap between the target ground state and the first excited states, in a similar fashion to adiabatic quantum computing (AQC). In AQC, the minimum instantaneous gap between the time-dependent ground state and first excited state classifies the hardness of the optimization problem as it determines the minimum time scale necessary for successful adiabatic state preparation [41]. In QAOA, instead, where continuous evolutions are replaced by digital steps and it is not possible to identify a continuous time-dependent generator of the quantum circuit, this role is played by the gap of the target Hamiltonian. This is a consequence of the fact that the variational procedure is more complicated for problem Hamiltonians with small gaps, as we are going to discuss later on.
@@ II.1 Fully-connected spin model
where $\sigma_{i}^{Z},\sigma_{j}^{Z}$ are the Pauli matrices acting on respectively the $i^{\text{th}}$ and $j^{\text{th}}$ spins and $J_{ij}=\mathcal{U}([-1,1])$ are random couplings distributed according to the uniform distribution $\mathcal{U}$ in the given range. The cost function of QAOA, as discussed in Sec. III, is given by the expectation value of $H_{T}$ on a parametric quantum circuit. As a proof of principle we focus on random instances with a small number of qubits, $N=5$ (in Appendix A, we show that our results also hold for larger systems), and analyze $n=600$ samples (which gives an accurate sampling of the coupling distribution without requiring unfeasible computational resources) of the disordered interactions, each of them with different couplings $J_{ij}$. In the case $J_{ij}=1$ finding the ground state of the Hamiltonian of Eq. (2) is the same as finding the maximum of the cost function $f$ defined in Eq. (1). Due to the $Z_{2}$ symmetry of the Hamiltonian of Eq. (2) all the eigenvalues are doubly degenerate.
@@ III QAOA and variants
QAOA is a hybrid quantum-classical algorithm proposed by Farhi et al [9]. It consists of a quantum part and a classical part. The objective of the algorithm is to find the ground state of a target Hamiltonian $H_{T}$. The algorithm works as follows. We consider two Hamiltonians: $H_{T}$, representing our target Hamiltonian whose ground state solves our optimization problem, and $H_{X}$, a one-body Hamiltonian easy to diagonalize, encoding quantum fluctuations:
@@
|$\ket{\psi_{\text{2CD}}^{(p)}(\vec{\beta},\vec{\gamma},\vec{\alpha},\vec{\delta},\vec{\zeta})}=$
@@ IV Results
In this section, we discuss results achieved by studying the disordered fully-connected spin model of Eq. (2) within the QAOA, QAOA-CD, and QAOA-2CD frameworks. The classical optimization is performed using the L-BFGS-B method [65]. For each algorithm (QAOA, QAOA-CD, QAOA-2CD) and at each step, we start the optimization with $20$ randomly selected sets of initial angles, and select the best result from these optimizations. While this approach does not guarantee a global optimum, the optimization can be refined by increasing the number of initial angle choices when needed. Our results are averaged over a set of $n=600$ random instances. Disorder-averaged quantities are calculated as
@@
The fidelity of Eq. (16) as well as the residual energy of Eq. (12) are used in the following to assess the performances of the algorithm and eventually connect them with the spectral properties of $H_{T}$. In quantum annealing [41, 54, 66], the minimal gap during the annealing dynamics determines the fidelity of the algorithm due to the occurrence of Landau-Zener transitions. In QAOA such strict relation between the gaps and the fidelity is not known nor well established, however a dense spectrum in the vicinity of the ground state, makes the system more prone to be “captured” by local minima during the hybrid optimization procedure [67]. To highlight the role of the gaps in the QAOA, in the following we will relate the QAOA performances for various instances keeping track of the distance between ground and the first excited state. In this paper we define this quantity as $\Delta_{eg}$.
@@
Figure 2: (a) Residual energy versus QAOA step $p$. Grey lines represent $\varepsilon_{\mathrm{res,}i}^{(p)}$ of all the $n=600$ random instances and the red line their average value $\braket{\varepsilon_{\mathrm{res}}}$. (b-d) Residual energy $\varepsilon_{\mathrm{res,}i}^{(p)}$ as a function of the minimal gap $\Delta_{eg}$ for QAOA at steps (b) $p=1$, (c) $p=2$ and (d) $p=7$. Three distinct regions based on the energy gap separating the ground state from the first excited state can be singled out for each instance: I-zone (green circles), corresponding to instances with $\Delta_{eg}\in\left[0,0.44\right]$, II-zone (blue triangles) for instances with $\Delta_{eg}\in\left[0.44,1.02\right]$, III-zone (red squares) for instances with $\Delta_{eg}\in\left[1.02,4.74\right]$. The horizontal line within each sector depicts the corresponding average value of that sector. We see that for $p=1$ these average values grow with increasing of $\Delta_{eg}$. This trend starts to change from step $p=2$ (panel c), whereas III-zone (panel d) shows a trend inversion (average value decreasing with the minimal gap).
@@
Figure 3: (a) Fidelity versus QAOA step $p$. Grey lines represent values $F_{i}^{(p)}$ of all $n=600$ random instances and the red line their average value $\braket{F}$. (b-d) Fidelity $F_{i}^{(p)}$ as a function of the minimal gap $\Delta_{eg}$ for QAOA at step (b) $p=1$, (c) $p=2$ and (d) $p=7$. In panels (b-d) the three distinct regions described in caption of Fig. 2 are evidenced. The horizontal line within each sector depicts the corresponding average value of that sector. For each step $p$ the fidelity closest to one is obtained for instances with $\Delta_{eg}$ belonging to III-zone.
@@ IV.1 QAOA
In Fig. 2(a) we show the residual energy as a function of the number $p$ of iterations of QAOA. Each grey line corresponds to the residual energy $\varepsilon_{\mathrm{res,}i}^{(p)}$ for a random instance, while the red bold curve indicates the average residual energy $\braket{\varepsilon_{\mathrm{res}}}$ over all the $n$ instances. The average residual energy at the step $p=10$ is of the order of $\braket{\varepsilon_{\mathrm{res}}^{(10)}}\simeq(2\pm 3)\cdot 10^{-3}$, consistent with a zero residual energy at large $p$ which means that, on average, QAOA gives a good approximation of the exact solution at $p=10$. However the finite standard deviation indicates that some of the instances have yet to converge to their respective ground states, justifying the need for counterdiabatic corrections at finite $p$.
Is it possible to identify a common pattern among the instances that converge more slowly? In order to answer this question, we investigate the connection between the gap $\Delta_{eg}$ and complexity of the algorithm, following what is done in Ref. [68] for quantum annealing. In Figs. 2(b–d), we show scatter plots of the residual energy of each instance as a function of its gap $\Delta_{eg}$, at some fixed steps $p$. In particular, we split the sorted gaps into three sets of equal cardinality (200 instances per set): in I there are problems with $\Delta_{eg}\in\left[0,0.44\right]$, in II problems with $\Delta_{eg}\in\left[0.44,1.02\right]$ and in III problems with $\Delta_{eg}\in\left[1.02,4.74\right]$. In panel (b) we show the scatter plots for $p=1$, in (c) those for $p=2$ and in (d) those for $p=7$. In panel (b), with $p=1$, we see that the larger is the gap the bigger is the residual energy; this behavior is confirmed by the average value of residual energy in each region. The reason is that, at the first steps of QAOA, the parameters are not sufficient to accurately describe the problem’s ground state. As a result, higher energy components of the QAOA variational guess over the computational basis states are detrimental for the algorithm. Starting from step $p=2$ we note that this trend reverses and becomes more and more marked as the number of steps $p$ increases, as we see comparing panels (c) and (d). In panel (c) the average residual energies of the regions II and III have similar values, then in the panel (d), for $p=7$, the average in the region III is lower than that in region I. The numerical values of the average residual energies in each region can be found in Appendix A.
@@
However, the correlation between $\Delta_{eg}$ and the values of residual energy and fidelity in QAOA could be related to the fact that in the presence of a dense spectrum of the Hamiltonian $H_{T}$ there is a large likelihood of identifying local minima in close proximity to the ground state energy. To better understand why small gaps influence the performance of QAOA, we analyzed the path followed by our optimizer routine for three instances characterized by different values of the spectral gap, one per each zone, at the first step of QAOA, where the parameter space is two-dimensional and can be visualized. Our analysis shows that in all three cases the minimizer can quickly find the global minimum regardless of the gap, but the variational ansatz itself is a worse approximation of the true ground state when the minimum gap is smaller. In order to better understand how the gap is important in the minimizer, the end of section IV.2 focuses on different roles played by the residual energy and the fidelity in finding the solution for QAOA and QAOA-2CD.
@@ IV.2 QAOA-CD and QAOA-2CD
Instances with $\Delta_{eg}\sim 0$ exhibit quasi-degeneracy between the ground and first excited states, leading to the QAOA state converging to a solution that reflects an overlap between these states [see Fig. 3(a)]. This behavior contrasts with the QAOA-2CD state, where, by step $p=7$, the fidelity reaches $1$ [see Fig. 4(d)]. From this observation, it appears that the role of the counterdiabatic correction in QAOA is analogous to that of the counterdiabatic potential in quantum annealing: both prevent transitions from the ground state to excited states, regardless of the system’s gap $\Delta_{eg}$. The key distinction is that in QA, gaps are dynamical, while in QAOA, $\Delta_{eg}$ pertains to the gap of the target Hamiltonian $H_{T}$.
@@
Figure 6: Fidelity $F_{i}^{(p)}$ of QAOA-2CD for all $n=$ 600 random instances versus $\Delta_{eg}$ for $p=1$ (a) and $p=2$ (b). Each point in the scatter plot represents the fidelity $F_{\mathrm{2CD,i}}^{(p)}$ for a single instance. Residual energy $\varepsilon_{\text{2CD,i}}^{(p)}$ for all $n=$ 600 different instances vs $\Delta_{eg}$ of QAOA-2CD for $p=1$ (c) and $p=2$ (d). Graphs delineate three distinct regions based on the energy gap $\Delta_{eg}$, as illustrated in caption of Fig. 2. The horizontal line within each sector depicts the corresponding average value of that sector. Fidelity in QAOA-2CD increases with $\Delta_{eg}$ as for QAOA. It is not possible to appreciate a specific behavior for residual energy.
@@
In Fig. 7 we examine three instances from distinct zones. In particular we focus on the instance with $\Delta_{eg}=0.0021$ from I-zone (green curve), which has the smallest $\Delta_{eg}$ among all the instances analyzed; instance with $\Delta_{eg}=1.9845$ from II-zone (blue curve); and instance with $\Delta_{eg}=4.7457$ from III-zone (orange curve), the instance with the largest value of $\Delta_{eg}$. In particular Fig. 7(a, b) shows the fidelity of these three instances with the ground state of $H_{T}$, |$\ket{\psi_{T}}$ for QAOA and QAOA-2CD respectively. In Fig. 7(c, d) we represent the overlap between QAOA and QAOA-2CD respectively and the first excited state.
By comparing panels (a) and (c), we observe that, at convergence, the instance with the smallest gap (orange curve) in QAOA exhibits an overlap between the ground and first excited states. This is because, with a very small $\Delta_{eg}$, the ground and first excited states are quasi-degenerate, so QAOA effectively finds the solution in the subspace spanned by both states. However, in panels (b) and (d), we see that counterdiabaticity can resolve this quasi-degeneracy: despite requiring more steps $p$ than other instances, the instance with the smallest gap in QAOA-2CD converges to the true ground state, despite the degeneracy.
(a)                          (b)
Refer to caption Refer to caption
(c)                          (d)
Refer to caption Refer to caption
Figure 7: (a) Fidelity between the QAOA state and the true ground state. (b) Excited-state probability of the optimal QAOA state. (c) Fidelity between the QAOA-2CD state and the true ground state. (d) Excited-state probability of the optimal QAOA-2CD state. Results are shown for three instances with different gap values $\Delta_{eg}$.
@@
Table A2: Average residual energies $\braket{\varepsilon_{\mathrm{res}}}$, $\braket{\varepsilon_{\mathrm{res,CD}}}$ and $\braket{\varepsilon_{\mathrm{res,2CD}}}$ for QAOA, QAOA-CD and QAOA-2CD, respectively, together with their respective standard deviations for all the analyzed $p$ steps. For each fixed $p$, the QAOA-2CD algorithm outperforms both the QAOA-CD and QAOA algorithms. Value denoted as $\sim 0$ have standard deviations $\leq 10^{-4}$.
@@ V Conclusions
To summarize, in this manuscript we have applied digitized counterdiabatic QAOA algorithms to study fully-connected spin models. In particolar we have compared the performances of QAOA, QAOA-CD, QAOA-2CD on a fully-connected disordered MaxCut problem with $N=5$ vertices. We analyzed $n=600$ random instances quantifying the quality of the algorithm with the residual energy (i. e., distance between the energy of the calculated solution and the exact ground-state energy) and the fidelity (i. e., the overlap between the exact ground state and the calculated state). Consistently with our prior findings [40], we observe that the QAOA-2CD overperforms the other variants, at the price of an increased number of classical parameters to optimize at fixed p. On the other hand, QAOA-CD and QAOA-2CD perform similarly to QAOA for a fixed number of variational parameters (not shown) [40].
We also analyzed the performances of the algorithms in connection with the gap between the ground state manifold and the first excited state, showing that in general, large gap instances are easier to be solved with this approach. While the role of spectral gaps are well known in QA, in QAOA their role is more subtle and this result is somehow nontrivial. The underlying idea, however, is that a dense spectrum in the vicinity of the ground state makes either the system more prone to be “captured” by local minima during the hybrid optimization procedure or, more likely, the overlap of the variational ground state with the excited states of $H_{T}$ not negligible. This approach could be used as a starting point to train a neural network with datasets sorted by spectral gaps to approximate efficiently the QAOA angles as done in Ref. [68] for QA.